OpenStudy (anonymous):

how do you multiply and divide radical 3 by an improper fraction without converting the fraction to a decimal? leave the radical in your answer.

OpenStudy (asnaseer):

dividing by a fraction is the same as multiplying by the reciprocal of the same fraction. i.e.:\[a\div \frac{b}{c}=a\times\frac{c}{b}\]this should help you answer your own question. let me know if you need more clarification.

OpenStudy (anonymous):

but what if i was multiplying an improper fraction like 14/3 and i wanted to multiply it by radical 3. how would i do that? you can keep the radical in your answer. Also how would i divide the radical by the improper fraction?

OpenStudy (asnaseer):

lets take an example:

OpenStudy (asnaseer):

suppose you wanted to divide 3 by 27/29, this is how you would do it:

OpenStudy (asnaseer):

\[3\div\frac{27}{29}=3\times\frac{29}{27}\]the 27 in the denominator can be divided by 3 to get 9, leaving us:\[=\frac{29}{9}\]does that help?

OpenStudy (asnaseer):